# The Adaptivity Barrier in Batched Nonparametric Bandits: Sharp Characterization of the Price of Unknown Margin

### v1

@@ 4 The RoBIN algorithm: optimal adaptation to unknown margin
We now introduce RoBIN (RObust batched algorithm with adaptive BINning) to tackle batched contextual bandits with unknown margin. The algorithm builds on the BaSEDB framework from [22], but introduces a new design principle: the batch schedule and split factors are selected based on the solution to the key variational convex program (8). As we shall see, this design principle enables robust adaptation across all margin parameter values.
@@ 4.1 The BaSEDB framework with fixed parameters
Next, we turn to the arm elimination part in Algorithm 2. The underlying idea is based on Successive Elimination (SE) from the bandit literature [11, 31, 15]. An arm is eliminated from the active arm set $\mathcal{I}_{C}$ if the revealed rewards from this batch provide sufficient evidence of the suboptimality of this arm. For any node $C\in\mathcal{T}$, denote by $m_{C,i}\coloneqq\sum_{t=t_{i-1}+1}^{t_{i}}\mathbf{1}\{X_{t}\in C\}$ the number of times the covariates go into the bin $C$ during the $i$-th batch. For $k\in\{1,-1\}$, define the empirical estimate arm $k$’s reward in bin $C$ during batch $i$ as
@@ 5.1 Proof overview
Let $\bm{u}^{\star}$ be the minimizer of the variational problem (8). For each $1\leq i\leq M$, let $\alpha_{i}\in\mathcal{K}$ be the margin parameter given by Proposition 3 such that
$\psi_{M}^{\star}=\;\eta_{i}\bigl(\bm{u}^{\star},\alpha_{i}\bigr)\;-h_{M}(\alpha_{i}).$ (14)
In particular, we have $\alpha_{i}\leq d/\beta$ for all $2\leq i\leq M$. Set $T_{i}=\lceil T^{u_{i}^{\star}}\rceil$ for $1\leq i\leq M-1$, and $T_{0}=0$, $T_{M}=T$.
We construct $M$ difficulty levels indexed by these $\alpha_{i}$:
  • Level 1: the arms are perfectly separated, and hence easy to distinguish;
  • Levels $i\geq 2$: Margin $\alpha_{i}$ is non-increasing, making identification of the optimal action more challenging. The level-$i$ instance requires $\approx T^{u_{i-1}^{\star}}$ samples to resolve.
@@
  • Hard-instance construction: We partition the covariate space into $M$ “stripes”, each operating at a different resolution. Within stripe $i$, we randomly place “active cells” with reward gaps aligned with margin $\alpha_{i}$. An algorithm cannot determine which cells are active without sufficient exploration.
  • Regret in the pivotal window: Let $\Gamma=\{0<t_{1}<t_{2}<\cdots<t_{M-1}<T\}$ be the (possibly adaptively chosen) grid points. For each $1\leq i\leq M$, define an event
    $A_{i}=\{t_{i-1}<T_{i-1}<T_{i}\leq t_{i}\}.$ (15)
    Note that $A_{1},\dots,A_{M}$ form a partition of the whole probability space. Roughly speaking, $A_{i}$ represents the event that the algorithm’s chosen grid points are suboptimal between the (i1)(i-1)-th and $i$-th batch. We show that if $A_{i}$ happens, the algorithm incurs large regret on level-$i$ cells within the window $T_{i-1},T_{i}]$.
  • Indistinguishability: We prove that observing $T_{i-1}$ samples cannot reliably distinguish between having active cells at level $i$ vs. level $i+1$. This forces the algorithm into a bad event for some level.
In what follows, we aim to construct $M$ mixture distributions $\{Q_{1},Q_{2},\ldots,Q_{M}\}$ such that $Q_{i}(A_{i})$ is large in at least one environment. We then argue that when $A_{i}$ happens, the regret inflation is necessarily large. The key challenge and complexity lie in the construction of $\{Q_{1},Q_{2},\ldots,Q_{M}\}$, which we describe in detail below.
@@ 5.2 The hard instance construction
$\xi_{m,j}(x):=\delta_{m}\,\phi^{\beta}\big(Mz_{m}\|x-q_{m,j}\|_{\infty}\big)\,\mathbf{1}{\{x\in C_{m,j}\}},\qquad\delta_{m}:=D_{\phi}\,(Mz_{m})^{-\beta},$ (17)
@@
$|S_{m}|=s_{m}:=\Big\lceil M^{-1}(Mz_{m})^{\,d-\alpha_{m}\beta}\Big\rceil,\quad 2\leq m\leq M,\quad|S_{1}|=s_{1}:=M^{d-1},$ (18)
@@ 5.4.1 An equivalent coin model
$N_{i}:=\sum_{t=1}^{N}\mathbf{1}\{I_{t}=i\},\quad R_{i}:=\sum_{t=1}^{N}\mathbf{1}\{I_{t}=i,Y_{t}=1\}.$
@@ 5.4.3 Proof of Lemma 5
Denote by $\bm{N}=(n_{1},\dots,n_{z})$ the multinominal vector which counts the number of times each coin is flipped. Conditioned on $\bm{N}$, coins are independent under both $H_{0}$ and $H_{1}$. Denote by $\bm{R}=(R_{1},\dots,R_{z})$. For the null model,
@@
$\mathbb{E}_{\sigma}\!\left[\prod_{i\in S}r_{i}(\sigma_{i})\right]=\prod_{i\in S}m_{i},\quad\text{where}\quad m_{i}:=\frac{1}{2}\big(r_{i}(+1)+r_{i}(-1)\big).$
@@
where the second step is due to the negative association of multinomial random variables [25], and the last step applies the PGF the multinomial distribution. We reach
@@
$\mathbb{E}_{P_{0}}[\mathbb{E}[\Lambda^{2}\mid\bm{N}]]$ $=\mathbb{E}_{S,S^{\prime}}\left[(g_{\delta/\sqrt{z}}(n))^{|S\cap S^{\prime}|}.\right]\leq\exp\left(\frac{s^{2}}{z}(g_{\delta/\sqrt{z}}(n)-1)\right).$ (27)
@@
6 Related work
Contextual bandits.
The concept of contextual bandits was introduced by [47]. For linear contextual bandits, [4, 1, 17, 7, 33] established regret guarantees in both low- and high-dimensional settings. Meanwhile, modeling the mean reward function as a smooth function of the contexts was studied in [48]. [40] proved a minimax regret lower bound for this setup and designed an upper-confidence-bound-type (UCB-type) policy to attain the near-optimal rate. [31] refined this result by proposing the Adaptively Binned Successive Elimination (ABSE) policy that can also adapt to the unknown margin parameter in the fully online setting. Additional insights in nonparametric contextual bandits were obtained in [34, 37, 18, 21, 42, 19, 8].
Margin condition in classification.
The margin condition originates from nonparametric classification, where it was studied by [3]. This condition, often referred to as the Tsybakov margin condition, quantifies how well-separated the optimal decision boundary is and directly governs learning rates. Its adaptation to contextual bandits was initiated by [16, 40, 31], who showed that the margin parameter $\alpha$ fundamentally shapes the complexity of bandit learning. In the online setting, adaptation to unknown $\alpha$ is possible without additional regret cost [31]. However, under batching constraints, as explored in our work, such adaptivity becomes costly, giving rise to a new barrier.
Batch learning.
The multi-armed bandit problem under the batched setting was studied by [32, 15]. Batch learning in linear contextual bandits was studied by [20, 39, 41] and [38, 45, 12] further considered the problem with high-dimensional covariates. [22, 2] studied the nonparametric contextual bandit problem under the batch constraint. [27, 26] developed batched Thompson sampling algorithms. [13] considered the Lipschitz continuum-armed bandit problem under the batched setting. Further insights in batched bandits were developed in [49, 23, 24, 28]. Another related topic is online learning with switching costs [10]. Best arm identification with limited rounds of interaction has been studied by [43]. Reinforcement learning with low switching costs has been considered by [5, 50, 14, 46, 36].
Adaptation.
In the fully online setting, adaptivity to the margin parameter is feasible at no extra cost [31]. One might ask whether the same is true for the smoothness parameter. The answer is negative: even without the batch constraint, adaptation to smoothness is impossible [29, 19, 9]. Minimax regret rates depend explicitly on the Hölder smoothness $\beta$, and no single procedure can achieve the optimal rate simultaneously across different values of $\beta$. This impossibility parallels classical results in nonparametric estimation and classification, where smoothness adaptation requires additional structure or necessarily incurs a penalty [35, 19, 8]. Thus, the margin parameter is the quantity of genuine interest: it admits free adaptation online, yet—as we shall show—becomes costly under batching.
@@ 7 Discussion
This work provides a complete characterization of the cost of adaptivity in batched nonparametric contextual bandits. By introducing regret inflation, we quantify how much additional regret is unavoidable when the margin parameter $\alpha$ is unknown. Our main finding is that this adaptivity cost scales as $T^{\psi^{*}_{M}}$, where $\psi^{*}_{M}$ is the value of a convex variational problem depending on the number of batches $M$, the smoothness $\beta$, and the dimension $d$. The matching upper and lower bounds show that this exponent captures the exact statistical price of limited adaptivity, up to logarithmic factors.
Phase transition in adaptivity.
A central insight of our analysis is the existence of a sharp threshold in the number of batches. When $M\gtrsim\log\log T$, the inflation exponent $\psi^{*}_{M}$ vanishes, implying that adaptivity to the unknown $\alpha$ is essentially free: a learner constrained to $O(\log\log T)$ updates can match the oracle rate as if $\alpha$ were known. Below this threshold, however, $\psi^{*}_{M}$ decreases geometrically in $M$, revealing a quantitative tradeoff between adaptivity and feedback granularity.
Algorithmic implications.
The constructive algorithm RoBIN demonstrates that optimal adaptivity can be achieved through simple scheduling guided by the variational program. The resulting batch schedule offers a practical prescription for designing batched exploration policies. In particular, it provides a practical rule for choosing the smallest $M$ that ensures near-oracle performance: select $M$ so that $T^{\psi^{*}_{M}}$ falls within the desired regret tolerance.
Limitations and future directions.
@@
Several natural extensions remain open. First, our analysis assumes two arms. Extending these results to multi-armed settings is certainly interesting. Second, when both the margin and smoothness parameters are unknown, one could ask whether a unified adaptive strategy is possible. Smoothness adaptation is known to require additional structure or to suffer unavoidable penalties even in the fully online case [35, 19, 8]. Batching further complicates this, because the size of the first batch needs to be specified before smoothness can be estimated. A joint variational formulation over $(\alpha,\beta)$ may shed light on this problem. Last but not least, the regret inflation framework may extend to other batched bandit models. For instance, in sparse linear contextual bandits [38], current algorithms rely on knowing an upper bound on the sparsity level. Quantifying the price of unknown sparsity or other complexity parameters would broaden the scope of our results.
@@ Acknowledgements
C.M. was partially supported by the National Science Foundation via grant DMS-2311127 and DMS CAREER Award 2443867.
Appendix A Minimax regret when margin is known
In this section, we prove Proposition 2.

### v2

@@ 4 The RoBIN algorithm: optimal adaptation to unknown margin
We now introduce RoBIN (RObust batched algorithm with adaptive BINning) to tackle batched contextual bandits with unknown margin. The algorithm builds on the BaSEDB framework from Jiang and Ma (2025), but introduces a new design principle: the batch schedule and split factors are selected based on the solution to the key variational convex program (8). As we shall see, this design principle enables robust adaptation across all margin parameter values.
@@ 4.1 The BaSEDB framework with fixed parameters
Next, we turn to the arm elimination part in Algorithm 2. The underlying idea is based on Successive Elimination (SE) from the bandit literature (Even-Dar et al., 2006; Perchet and Rigollet, 2013; Gao et al., 2019). An arm is eliminated from the active arm set $\mathcal{I}_{C}$ if the revealed rewards from this batch provide sufficient evidence of the suboptimality of this arm. For any node $C\in\mathcal{T}$, denote by $m_{C,i}\coloneqq\sum_{t=t_{i-1}+1}^{t_{i}}\mathbf{1}\{X_{t}\in C\}$ the number of times the covariates go into the bin $C$ during the $i$-th batch. For $k\in\{1,-1\}$, define the empirical estimate arm $k$’s reward in bin $C$ during batch $i$ as
@@ 5.1 Proof overview
Let $\bm{u}^{\star}$ be a minimizer of the variational problem (8). By Proposition 3(ii), there exist margin parameters $\alpha_{1}=\infty\geq\alpha_{2}\geq\cdots\geq\alpha_{M}$ with $\alpha_{i}\leq d/\beta$ for $i\geq 2$ such that
$\psi_{M}^{\star}\;=\;\eta_{i}(\bm{u}^{\star},\alpha_{i})-h_{M}(\alpha_{i})\quad\text{for each }i\in[M].$ (14)
Define batch cutoffs $T_{0}=0$, $T_{i}=\lceil T^{u_{i}^{\star}}\rceil$ for $1\leq i\leq M-1$, and $T_{M}=T$. We interpret index $i$ as a difficulty level: level 1 has a constant gap (easy), while levels $i\geq 2$ have nonincreasing margin parameters (harder). Crucially, resolving level $i$ requires on the order of $T^{u_{i-1}^{\star}}$ samples, which is exactly the size of the preceding window.
@@
  • Hard-instance construction (Section 5.2). We partition the covariate space into $M$ “stripes”, each operating at a different resolution. Within stripe $i$, we randomly place “active cells” with reward gaps aligned with margin $\alpha_{i}$. An algorithm cannot determine which cells are active without sufficient exploration.
  • Regret in the pivotal window (Section 5.3). Let $\Gamma=\{0<t_{1}<t_{2}<\cdots<t_{M-1}<T\}$ be the (possibly adaptively chosen) grid points by the policy. Define the bad events
    $A_{i}=\{t_{i-1}<T_{i-1}<T_{i}\leq t_{i}\},\qquad i=1,\ldots,M,$ (15)
    which form a partition of the sample space. Roughly speaking, $A_{i}$ represents the event that the algorithm’s chosen grid points are suboptimal between the (i1)(i-1)-th and $i$-th batch. We show that if $A_{i}$ happens, the algorithm incurs large regret on level-$i$ cells within the window $T_{i-1},T_{i}]$.
  • Indistinguishability across consecutive levels (Section 5.4). We define mixtures $Q_{i}$ by randomizing the active sets and signs for levels $m\leq i$. The key is that observations up to time $T_{i-1}$ cannot reliably distinguish whether the environment is drawn from $Q_{i}$ or $Q_{i+1}$:
[…]
    This is proved via an equivalent coin model (biased vs. fair coins) and a $\chi^{2}$ bound (Lemma 5), followed by Pinsker. By a telescoping argument over $i$ and the triangle inequality, there exists an index $i^{\star}$ such that
[…]
    Thus, on at least one level, the pivotal window occurs with nontrivial probability under the corresponding mixture.
@@ 5.2 The hard instance construction
$\xi_{m,j}(x)\coloneqq\delta_{m}\,\phi^{\beta}\big(Mz_{m}\|x-q_{m,j}\|_{\infty}\big)\,\mathbf{1}{\{x\in C_{m,j}\}},\qquad\delta_{m}\coloneqq D_{\phi}\,(Mz_{m})^{-\beta},$ (17)
@@
$|S_{m}|=s_{m}\coloneqq\Big\lceil M^{-1}(Mz_{m})^{\,d-\alpha_{m}\beta}\Big\rceil,\quad 2\leq m\leq M,\quad|S_{1}|=s_{1}\coloneqq M^{d-1},$ (18)
@@ 5.4.1 An equivalent coin model
$N_{i}\coloneqq\sum_{t=1}^{N}\mathbf{1}\{I_{t}=i\},\quad R_{i}\coloneqq\sum_{t=1}^{N}\mathbf{1}\{I_{t}=i,Y_{t}=1\}.$
@@ 5.4.3 Proof of Lemma 5
We bound the chi-squared divergence using the second-moment representation, i.e.,
$\chi^{2}(Q,P_{0})=\mathbb{E}_{P_{0}}[\mathbb{E}[\Lambda^{2}\mid\bm{N}]]-1,$
where we further condition on $\bm{N}=(N_{1},\dots,N_{z})$, the multinominal vector which counts the number of times each coin is flipped.
@@
Conditioned on $\bm{N}$, coins are independent under both $H_{0}$ and $H_{1}$. Denote by $\bm{R}=(R_{1},\dots,R_{z})$. For the null model,
@@
$\mathbb{E}_{\sigma}\!\left[\prod_{i\in S}r_{i}(\sigma_{i})\right]=\prod_{i\in S}m_{i},\quad\text{where}\quad m_{i}\coloneqq\frac{1}{2}\big(r_{i}(+1)+r_{i}(-1)\big).$
@@
where the second step is due to the negative association of multinomial random variables (Joag-Dev and Proschan, 1983), and the last step applies the probability generating function of the binomial distribution. We reach
@@
$\mathbb{E}_{P_{0}}[\mathbb{E}[\Lambda^{2}\mid\bm{N}]]$ $=\mathbb{E}_{S,S^{\prime}}\left[(g_{\delta/\sqrt{z}}(n))^{|S\cap S^{\prime}|}\right]\leq\exp\left(\frac{s^{2}}{z}(g_{\delta/\sqrt{z}}(n)-1)\right).$ (27)
@@ 6 Discussion
This work provides a complete characterization of the cost of adaptivity in batched nonparametric contextual bandits. By introducing the notion of regret inflation, we quantify how much additional regret is unavoidable when the margin parameter $\alpha$ is unknown. Our main finding is that this adaptivity cost scales as $T^{\psi_{M}^{\star}}$, where $\psi_{M}^{\star}$ is the value of a convex variational problem depending on the number of batches $M$, the smoothness $\beta$, and the dimension $d$. The matching upper and lower bounds show that this exponent captures the exact statistical price of limited adaptivity, up to logarithmic factors.
6.1 Phase transition in batched contextual bandits
Our results, together with those in Jiang and Ma (2025) (see also Proposition 2 in this paper), reveal a sharp phase transition in the interplay between batching, adaptivity, and oracle knowledge.
Figure 4 provides a clear visual summary of this landscape. The green curve represents the normalized regret exponent of an oracle algorithm (e.g., BaSEDB) that knows $\alpha$, while the red curve shows the regret exponent of the optimal adaptive algorithm (e.g., RoBIN). The gap between them, which we characterize by $\psi_{M}^{\star}$, is precisely the “price of adaptation.” Our analysis explains the behavior of this gap, revealing three qualitatively distinct regimes determined by the number of batches $M$ relative to the horizon $T$:
  1. 1.
    When $M\leq c\log\log T$ for a sufficiently small constant $c$, the batch constraint itself dominates. In this under-batched regime, even an oracle algorithm that knows the true margin parameter $\alpha$ cannot attain the fully online regret rate.
  2. 2.
    In the intermediate range $c\log\log T<M<C\log\log T$, knowing $\alpha$ becomes beneficial. An oracle algorithm with knowledge of the true margin parameter can achieve the online rate up to logarithmic factors, but ignorance of $\alpha$ induces regret inflation quantified by $\psi_{M}^{\star}$. This transition regime therefore marks the true adaptivity barrier, where batching and parameter uncertainty interact nontrivially.
  3. 3.
    Finally, when $M\geq C\log\log T$ for a sufficiently large constant $C$, the regret inflation becomes constant. In this fully adaptive regime, the proposed algorithm RoBIN matches the performance of the online oracle up to polylogarithmic factors, and adaptation to the unknown $\alpha$ becomes effectively free.
Figure 4: Phase diagram illustrating the three regimes of adaptivity under the batch constraint. The $x$-axis represents the number of batches $M$, and the $y$-axis shows normalized regret. The green curve corresponds to the performance of the optimal algorithm that knows the margin parameter $\alpha$, while the red curve corresponds to the optimal adaptive algorithm without knowledge of $\alpha$. The dotted line shows the exponent for an optimal online algorithm.
These three regimes collectively characterize the adaptivity landscape of batched contextual bandits: batching imposes a fundamental statistical cost only below the doubly logarithmic threshold, beyond which adaptation incurs no additional penalty. (More precisely, the lower and upper thresholds satisfy $c\asymp 1/\log(1/\gamma(\alpha))$ and $C\asymp 1/\log(1/\gamma_{\mathrm{max}})$, where $\gamma_{\mathrm{max}}=\gamma(d/\beta)$ and $\gamma(\alpha)\leq\gamma_{\mathrm{max}}$, giving the ordering $c<C$.)
6.2 Extension to general knowledge sets
In practice, one may have partial prior information about the possible range of the margin parameter. This can be represented by restricting the knowledge set $\mathcal{K}$ to a subset of $0,d/\beta]\cup\{\infty\}$, as in Definition 1. Our main results, in particular Theorem 1, continue to hold whenever $\mathcal{K}\subseteq[d/(2\beta),d/\beta]\cup\{\infty\}$.
Theorem 3 (General knowledge set).
Fix $\mathcal{K}$ to be a compact subset of $d/(2\beta),d/\beta]\cup\{\infty\}$. Let $\psi_{M,\mathcal{K}}^{\star}$ denote the optimal value of the variational problem defined in (8) with knowledge set $\mathcal{K}$. Then there exist constants $c_{1},c_{2}>0$, independent of $T$ and $M$, such that
$c_{1}M^{-8}(\log T)^{-1}T^{\psi_{M,\mathcal{K}}^{\star}}\;\leq\;\inf_{\Gamma,\pi}\mathrm{RI}(\Gamma,\pi)\;\leq\;c_{2}\,M^{5}(\log T)\,T^{\psi_{M,\mathcal{K}}^{\star}}.$
The proof follows closely that of Theorem 1, except that Lemma 13 now holds automatically. Indeed, for all $\alpha\in[d/(2\beta),d/\beta]\cup\{\infty\}$, we have $\gamma(\alpha)=\frac{(\alpha+1)\beta}{d+2\beta}\geq 1/2$, which guarantees that the successive hard-instance families $\{Q_{1},\ldots,Q_{M}\}$ cannot be reliably distinguished within a finite number of batches. This property—anchored at the boundary point $d/(2\beta)$—preserves the indistinguishability argument used in the lower bound proof. By contrast, when the knowledge set $\mathcal{K}$ extends below this threshold, $\gamma(\alpha)$ may fall below $1/2$, breaking the reduction and invalidating the argument.
Nevertheless, if one restricts the $M$-batch policy $(\Gamma,\pi)$ to use a fixed grid, that is, grid points $(t_{1},\ldots,t_{M-1})$ are chosen in advance, then $\psi_{M,\mathcal{K}}^{\star}$ still characterizes the optimal regret inflation over this restricted policy class.
Theorem 4 (General knowledge set, fixed grid).
Fix $\mathcal{K}$ to be a compact subset of $0,d/\beta]\cup\{\infty\}$. Then there exist constants $c_{1},c_{2}>0$, independent of $T$ and $M$, such that
$c_{1}M^{-8}(\log T)^{-1}T^{\psi_{M,\mathcal{K}}^{\star}}\;\leq\;\inf_{\Gamma,\pi}\mathrm{RI}(\Gamma,\pi)\;\leq\;c_{2}\,M^{5}(\log T)\,T^{\psi_{M,\mathcal{K}}^{\star}},$
where the infimum is taken over all $M$-batch policies with a fixed grid.
The upper bound follows directly from the performance guarantee of RoBIN. For the lower bound, since the policy class is restricted to fixed grids, the indistinguishability argument is unnecessary—one can directly choose the worst-case family $Q_{i}$ for the bad event $A_{i}$ (see Equation (15)). We omit the detailed proof for brevity.
6.3 Monotonicity of $\psi_{M}^{\star}$ in $M$
Empirically, $\psi_{M}^{\star}$ decreases monotonically with the number of batches $M$ (see Section 3.5). This behavior is not immediate from the definition of the variational objective in (8): while adding batches allows a finer schedule and hence lower regret, the benchmark $h_{M}(\alpha)$ also shrinks, making the improvement subtle to analyze. For small $M$, we can establish monotonicity formally:
Proposition 7.
For $M\leq 4$, we have $\psi_{M+1}^{\star}<\psi_{M}^{\star}$.
The proof appears in Section B.2.3. Whether this monotonicity extends to all $M$ remains an interesting open question.
6.4 Future directions
@@
Several natural extensions remain open. First, our analysis focuses on the two-armed setting; extending these results to multiple arms is a natural next step. Second, when both the margin and smoothness parameters are unknown, one may ask whether a unified adaptive strategy is possible. Smoothness adaptation is known to require additional structure or to incur unavoidable penalties even in the fully online case (Qian and Yang, 2016b; Gur et al., 2022; Cai et al., 2024). Under batch constraints, the challenge intensifies because the size of the initial batch must be determined before smoothness can be estimated. A joint variational formulation over $(\alpha,\beta)$ may shed light on this problem.
Finally, the regret inflation framework may extend to other batched bandit models. For example, in high-dimensional sparse linear contextual bandits (Ren and Zhou, 2023), existing algorithms rely on knowing an upper bound on the sparsity level. Quantifying the price of unknown sparsity or other structural complexity parameters would further broaden the reach of our theory.
@@ Acknowledgements
C.M. was partially supported by the National Science Foundation via grant DMS-2311127 and the CAREER Award DMS-2443867.
Appendix A Proof of oracle regret bounds (Proposition 2)
We split the proof into two cases: $\alpha=\infty$ and $\alpha\leq d/\beta$.